From GlueXWiki
Jump to: navigation, search

February 5, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Gas puzzle
    • gain for different wire diameters [2]
  2. CDC update (Mike)
  3. FDC update (Lubomir, Vlad)
  4. Electronics (Fernando, Cody)
  5. Engineering (Dave)
  6. Other


Participants: Curtis, Naomi (CMU), Beni, Simon, Eugene, Dave, Fernando, Chris, Nick, Vlad, and Lubomir (JLab).

Gas puzzle

- CO2 sensor sent to the company for repair; will have it next week. Initially, it will be connected to a laptop (serial to USB) with the GASLAB software. For the spring run we want to connect it to a terminal server, but then we need software to put the information in epics (Hovanes).

- Using Garfield, Lubomir compared the gains in the CDC with the two wire diameters, 17.8 um used in the prototype and 20 um used in the CDC. Plot of the response of simulated 55Fe source is linked above - about 70% increase in the gain from 20 to 17.8 um wire, when using Ar/CO2/propanol 58/42/0.33 gas mixture. Naomi: with the standard gas 50/50 they have seen just 10% increase. Lubomir will do simulations with the standard gas, as well.

- Gain measurements with 55Fe source will be done with both, the FDC spare package (standard 40/60 gas) and the FDC in the Hall with whatever gas is there, to make sure we understand the gas mixtures and the Garfield simulations. The spare package in EEL126 is being flushed with gas since last Friday. Beni is recovering the DAQ we used in 126, but there are some issues yet related to the new firmware.

CDC update

- Mike is working on the CDC efficiency. CDC is ready for taking data with the BCAL cosmic trigger, hopefully this weekend.

FDC update

- We discussed again the structures seen in the difference between the cathode reconstructed and the nominal wire position as a function of x and y - 2d plots and x-projections for all the cells in package 3 are attached. Note, there's a shift by 2mm for better 2d visualization. It is amazing, you see the same structure in all the cells, the projections can be fitted with almost the same sine function with an amplitude of ~200 microns. So far, no idea what is it.

- Vlad recalled that 2.5 years ago he put O-ring/cathode foil samples in the old Simon's prototype, to study the corrosion. The O-rings were pressed against the foils in the chamber grooves and the prototype was stored in EEL126. From the pictures linked above one can see that the sample with the EPDM O-ring has dark spots and some damage of the copper layer, while the sample with Viton O-rings (that we used in the FDC) looks almost unchanged.


- Out of 8 channels left in the FDC that have problems outside of the detector, Nick found 5 with problems in the 3M connectors. These have been fixed but not tested yet.